A person has refractive hypermetropia with a near point at 3 m. They are to be prescribed a 

set of glasses that will enable them to have normal close-in vision and the person’s glasses will 

typically sit 2 cm in front of their eyes. (a) What is the maximum optical power of this person’s 

eyes? (b) What is the optical power of the glasses required?


1
Expert's answer
2021-06-06T18:33:01-0400

a)

"D_{max}=\frac1f+\frac 1d=\frac 13+\frac{1}{0.02}=50.3~dpt,"

b)

"d_0=0.25~m,"

"D=\frac 1f+\frac{1}{d_0}=\frac 13+\frac{1}{0.25}=4.3~dpt."